# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ Abstract
Advances in architectural design, data availability, and compute have driven remarkable progress in semantic segmentation. Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making. The resulting reliance on point estimates has fueled interest in probabilistic segmentation, but the literature remains fragmented. In response, this review consolidates and contextualizes foundational concepts in uncertainty modeling, including the non-trivial task of distinguishing between epistemic and aleatoric uncertainty and examining their roles across four key downstream segmentation tasks, highlighting Active Learning as particularly promising. By unifying theory, terminology, and applications, we provide a coherent foundation for researchers and identify critical challenges, such as strong assumptions in spatial aggregation, lack of standardized benchmarks, and pitfalls in current uncertainty quantification methods. We identify trends such as the adoption of contemporary generative models, driven by advances in the broader field of generative modeling, with segmentation-specific innovation primarily in the conditioning mechanisms. Moreover, we observe growing interest in distribution- and sampling-free approaches to uncertainty estimation. We further propose directions for advancing uncertainty-aware segmentation in deep learning, including pragmatic strategies for disentangling different sources of uncertainty, novel uncertainty modeling approaches and improved Transformer-based backbones. In this way, we aim to support the development of more reliable, efficient, and interpretable segmentation models that effectively incorporate uncertainty into real-world applications.
@@ 1 Introduction
Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest (Szeliski, 2010). The advent of convolutional neural networks (CNNs) has led to major breakthroughs in this domain, with deep learning-based methods achieving state-of-the-art performance on large-scale datasets (Ronneberger et al., 2015; Shelhamer et al., 2014; Badrinarayanan et al., 2015), obtaining impressive scores with large-scale segmentation datasets (Lin et al., 2014; Cordts et al., 2016b; Richter et al., 2016). However, these models typically rely on strong assumptions and significant relaxations of the Bayesian learning paradigm, neglecting the uncertainty associated with their predictions. This lack of uncertainty modeling reduces both the reliability and interpretability of the predictions. In high-stakes applications, such as autonomous driving or medical diagnosis, this can have severe consequences. For instance, misclassifying adjacent objects in autonomous driving or overlooking uncertainty in lesion classification can both lead to critical decision-making errors.
@@
Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation, especially as interpretability and reliability have become central in data-driven applications. Extensive efforts have been made to align neural network optimization with Bayesian machine learning (Blundell et al., 2015; Guo et al., 2017; Kingma & Welling, 2013; Kendall & Gal, 2017), such as learning parameter distributions rather than point estimates to capture epistemic uncertainty. Additionally, explicitly modeling the likelihood of outputs enables estimation of aleatoric uncertainty. However, in practice, these two forms of uncertainty are often entangled or even conflicting (Mucsányi et al., 2024; de Jong et al., 2024; Valdenegro-Toro & Mori, 2022; Hüllermeier & Waegeman, 2021), making the distinction difficult to define and apply. This conceptual ambiguity presents significant barriers to understanding and implementation (Kahl et al., 2024; Der Kiureghian & Ditlevsen, 2009; Kirchhof et al., 2025). At the same time, uncertainty is often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective. As a result, many works lack rigorous theoretical grounding, focusing instead on empirical gains without clear definitions or principled formulations of uncertainty. Furthermore, many approaches are tailored to specific datasets or modalities, limiting generalizability. Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, and can be insufficiently contextualized within the domain of segmentation. This disconnect has resulted in a fragmented body of work, with inconsistent terminology, evaluation metrics, and assumptions. The resulting abundance of literature can be overwhelming, even for experienced researchers.
This paper aims to explore and clarify these concepts in the context of segmentation architectures and their impact on downstream tasks. Given these challenges and the renewed interest in uncertainty (Papamarkou et al., 2024; Kirchhof et al., 2025), we note that a comprehensive overview in the field remains limited. Due to its application-driven nature, most existing surveys adopt a medical perspective (Jungo & Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023). However, the connection between theoretical foundations and their diverse applications remains underexplored. Additionally, the study by Kahl et al. (2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks. In contrast, our work provides a comprehensive overview that spans the full range of established uncertainty quantification methods, unifies and contextualizes the underlying theory, and standardizes terminology and notation. This broader perspective helps clarify conceptual foundations and highlights cross-cutting insights across segmentation architectures and application domains. By the end of this paper, readers will have a clear understanding of the various forms of uncertainty, their relevance to segmentation tasks, and a comprehensive grasp of the key challenges and open research directions in the field.
[…]
@@
This review paper is structured as follows. Past work with significant impact in general image segmentation is presented in Section 2. Then, the theoretical framework and notation that govern the remainder of the paper are introduced in Section 3. The theory enables us to classify the approaches into two distinct but often related methods. That is, methods targeting either aleatoric or epistemic uncertainty, which can be distinguished by either modeling the feature or parameter distribution(s). The role of these concepts in image segmentation are treated in Sections 4 and 5, respectively, which include all architectures and approaches with significant impact on the field. We then correlate these modeling methods to specific applications in Section 6. Here, we also observe for which domains uncertainty quantification has been exploited. Then, this overview is discussed further in Section 7. This section is particularly valuable for researchers seeking to identify the most suitable model for their specific segmentation task. To encourage further research in this field, we provide future recommendations by highlighting key challenges, gaps and pitfalls in literature, concluding our review in Section 8. Figure 1 illustrates a brief overview of the sections.
@@ 2 Background
A common approach in image understanding involves labeling pixels according to semantic categories, which is the core of semantic segmentation. This technique is particularly well-suited for amorphous or uncountable subjects. In contrast, instance segmentation not only assigns class labels, but also distinguishes and delineates individual object occurrences. This makes it more appropriate for scenarios involving countable entities. A third variant, panoptic segmentation, unifies both semantic and instance-level labeling, offering a comprehensive view of scene composition. As summarized by Minaee et al. (2020), semantic segmentation has been performed using methods such as thresholding (Otsu, 1979), histogram-based bundling, region-growing (Dhanachandra et al., 2015), k-means clustering (Nock & Nielsen, 2004), watershedding (Najman & Schmitt, 1994), to more advanced algorithms such as active contours (Kass et al., 2004), graph cuts (Boykov et al., 2001), conditional and Markov random fields (Plath et al., 2009), and sparsity-based methods (Starck et al., 2005; Minaee & Wang, 2017). While the literature on segmentation is vast and rapidly evolving, a selection of backbone architectures has been particularly influential in shaping current probabilistic segmentation models. The focus here is not on exhaustiveness, but on those models most relevant to the development of uncertainty-aware approaches.
In particular, following the successful application of CNNs (LeCun et al., 1998), image segmentation experienced rapid progress driven by increasingly powerful and specialized deep architectures. Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan & Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 (Chen et al., 2017), and the MobileNetv3 (Howard et al., 2019) have also been commonly used. As the research progressed, increasing success has been observed with encoder-decoder models (Noh et al., 2015; Badrinarayanan et al., 2015; Yuan et al., 2019; Ronneberger et al., 2015). Initially developed for the medical applications, Ronneberger Ronneberger et al. (2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for many semantic segmentation and even general image generation architectures (Ho et al., 2020; Song et al., 2020), particularly in the medical domain. In fact, reports of recent research indicate that the relatively simple U-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
@@ 3 Probabilistic Image Segmentation
Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\*Y,\*X}$ that take values in $\mathcal{Y}\in\mathbb{Z}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then instance $\mathbf{y}$ can be considered as the ground-truth of a $K$-class segmentation task and instance $\mathbf{x}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively. Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
assuming independence of $\mathbf{x}$ from $\bm{\theta}$, and where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence). After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution from a new datapoint $\mathbf{x}^{*}$ can be denoted as
@@
As evident, both the variability in the empirical data and the inferred parameters of the model influence the predictive distribution. Hence, uncertainties stemming from the conditional likelihood distribution are classified as either aleatoric, implying from the statistical diversity in the data, or epistemic, which stems from the posterior, i.e. the variance of the model parameters. Encapsulating a particular uncertainty can therefore be achieved by incorporating parametrized stochasticity at either feature- or parameter-level. A straightforward approach to quantify any of these uncertainties is achieved by obtaining the predictive entropy, $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]$, i.e. the entropy of the predictive distribution $p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})$. Furthermore, disentangling these uncertainties can be achieved through its decomposition
@@
where $I$ represents the mutual information. In this way, it is theoretically possible to identify whether high-entropy predictions are due to model ignorance or statistical ambiguity inherent in the data generating process. Nonetheless, determining the nature of uncertainty is not often straightforward. For example, Hüllermeier & Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred". This sentiment is also shared by Der Kiureghian & Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic". Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic as we treat the realm of quantifying spatially correlated uncertainty. Especially with increasingly complex methodologies, treating the uncertainties as separate concepts is mostly theoretical (often even more philosophical) and highly non-trivial in practice.
Figure 2: Aleatoric uncertainty quantification by modeling pixel-level outputs as parameters of a Probability Mass Function.
@@ 3.1 Conventional segmentation
Regardless of the elegantly formulated Bayesian posterior, most practical approaches make use of so-called “deterministic” segmentation networks, which are trained by Maximum Likelihood Estimation (MLE) and is specified as
@@
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the model parameter variances approach zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution but can be achieved through Maximum A Posteriori (MAP) estimation with
@@
For example, assuming Gaussian or Laplacian priors leads to regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007). To model $p(\mathbf{y}|\mathbf{x},\bm{\theta})$, we make use of function $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ that infers the parameters of a Probability Mass Function (PMF). For instance, consider spatial image dimensions $D$ and a CNN with $\mathbf{a}=f_{\theta}(\mathbf{x})$. Then, we can write
@@
with channel-wise indexing over the denominator, which is commonly known as the SoftMax activation (Figure 2). This approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature. In fact, the approximated distribution can represent and localize uncertain regions. However, the implicit pixel-independence assumption
@@
omits information on structural variation in the segmentation masks. In conventional classification, factorizing the categorical distribution is typically regarded as a logical simplification. In probabilistic segmentation however, this assumption has caused the emergence of a distinct research direction (See Figure 3). The challenge of sampling of spatially coherent segmentation masks can be addressed either from the aleatoric or epistemic perspective, usually indicating reducability of the assumed uncertainty source. Both approaches have specific use cases, and each modeling choice comes with distinct advantages and limitations, which will be discussed in the following sections.
@@
Figure 3: Illustration of the likelihood function in segmentation models. Color intensities reflect the normalized confidence values that can be interpreted as probabilities. (a) The continuous output results in a horizontal gradient. (b) Maximum likelihood thresholding can be applied. (c) However, the coherence of the segmentation suffers when sampling.

### v8

@@ Abstract
Advances in architectural design, data availability, and compute have driven remarkable progress in semantic segmentation. Yet, these models often rely on relaxed Bayesian assumptions, omitting critical uncertainty information needed for robust decision-making. Despite growing interest in probabilistic segmentation to address point-estimate limitations, the research landscape remains fragmented. In response, this review synthesizes foundational concepts in uncertainty modeling, analyzing how feature- and parameter-distribution modeling impact four key segmentation tasks: Observer Variability, Active Learning, Model Introspection, and Model Generalization. Our work establishes a common framework by standardizing theory, notation, and terminology, thereby bridging the gap between method developers, task specialists, and applied researchers. We then discuss critical challenges, including the nuanced distinction between uncertainty types, strong assumptions in spatial aggregation, the lack of standardized benchmarks, and pitfalls in current quantification methods. We identify promising avenues for future research, such as uncertainty-aware active learning, data-driven benchmarks, transformer-based models, and novel techniques to move from simple segmentation problems to uncertainty in holistic scene understanding. Based on our analysis, we offer practical guidelines for researchers on method selection, evaluation, reproducibility, and meaningful uncertainty estimation. Ultimately, our goal is to facilitate the development of more reliable, efficient, and interpretable segmentation models that can be confidently deployed in real-world applications.
@@ 1 Introduction
Image segmentation entails pixel-wise classification of data, effectively delineating objects and regions of interest (Szeliski, 2010). The advent of convolutional neural networks (CNNs) has led to major breakthroughs in this domain (Ronneberger et al., 2015; Shelhamer et al., 2014; Badrinarayanan et al., 2015), obtaining impressive scores with large-scale segmentation datasets (Lin et al., 2014; Cordts et al., 2016b; Richter et al., 2016). However, these models typically rely on strong assumptions and significant relaxations of the Bayesian learning paradigm, neglecting the uncertainty associated within their predictions. This lack of uncertainty modeling reduces both the reliability and interpretability of these predictions. In high-stake applications, such as autonomous driving or medical diagnosis, this can have severe consequences. For instance, misclassifying objects in autonomous driving or overlooking uncertainty in lesion classification can both lead to critical decision-making errors.
@@
Fortunately, the merits of uncertainty quantification have been well-recognized in the field of CNN-based segmentation, especially as interpretability and reliability have become central in data-driven applications. Hence, extensive efforts have been made to align neural network optimization with Bayesian machine learning, such as learning parameter distributions rather than point estimates or to explicitly model the likelihood distribution of model features (Blundell et al., 2015; Guo et al., 2017; Kingma and Welling, 2013; Kendall and Gal, 2017). At the same time, uncertainty in image segmentation often treated as an auxiliary tool to enhance downstream performance, rather than as a primary modeling objective. Consequently, many studies lack a rigorous theoretical foundation, prioritizing empirical gains over principled formulations and clear definitions of uncertainty. This narrow focus frequently results in methods tailored to specific datasets or modalities, which limits their generalizability. Due to this application-driven nature, most existing surveys adopt a medical perspective (Jungo and Reyes, 2019; Kwon et al., 2020), often focused on specific modalities (McCrindle et al., 2021; Jungo et al., 2020; Ng et al., 2022; Roshanzamir et al., 2023). Additionally, the study by Kahl et al. (2024) introduces a valuable framework for benchmarking uncertainty disentanglement in semantic segmentation, but focuses primarily on empirical evaluation within specific tasks. Moreover, theoretical contributions often originate from adjacent fields like Bayesian deep learning or information theory, and can be insufficiently contextualized within the domain of segmentation. This disconnect has resulted in a fragmented body of work, with inconsistent terminology, evaluation metrics, and assumptions. The resulting abundance of literature can be overwhelming, even for experienced researchers.
Given the fragmented nature of the existing literature and the renewed interest in uncertainty (Papamarkou et al., 2024; Kirchhof et al., April 28, 2025), we note that a comprehensive overview in the field remains limited. Our central contribution is to unify core uncertainty modeling concepts across methods, tasks, and applications, thereby offering substantial theoretical insights and demonstrating their relevance across diverse domains. We unify and contextualize the underlying theory, and standardize terminology and notation. We systematically compare methodologies, link them to specific tasks, and connect those tasks to practical applications. Hence, this review is not only a conventional "zero-to-hero" survey; rather, it additionally serves as a cross-domain synthesis that bridges method developers, task specialists, and application-focused researchers. Our aim is to provide a cohesive framework for navigating and integrating the complex uncertainty modeling landscape. This broader perspective helps clarify conceptual foundations, highlight cross-cutting insights and provide a comprehensive grasp of the key challenges and open research directions in the field. In the absence of broad quantitative consensus, we adopt a robust, evidence-based methodology for making recommendations. By leveraging a deep qualitative analysis of the literature, we derive reliable method comparisons and guidelines that offer essential and practical guidance. However, this reliance on qualitative synthesis rather than quantitative evidence is an important limitation and our recommendations should be interpreted accordingly.
We begin by establishing the background on image segmentation in Section 2. Then, Section 3 introduces the notation and general theoretical principles of probabilistic image segmentation. Subsequently, we organize the review based on the source of introduced stochasticity: the feature level (Section 4) and the parameter level (Section 5). This structure aligns with the underlying theoretical decomposition of uncertainty and the utility of the modeled uncertainty for the downstream tasks and application domains discussed in Section 6. The review concludes by moving to the discussion (Section 7), where we highlight important pitfalls, challenges and opportunities in the field. This is followed by recommendations and guidelines for researchers (Section 8), before concluding in Section 9. The core structure and organizational flow of this paper is visualized in Figure 1, which serves as a top-level roadmap for the reader.
{forest}
@@ 2 Background
A common approach in image understanding involves labeling pixels according to semantic categories, which is the core of semantic segmentation. This technique is particularly well-suited for amorphous or uncountable subjects. In contrast, instance segmentation not only assigns class labels, but also distinguishes and delineates individual object occurrences. This makes it more appropriate for scenarios involving countable entities. A third variant, panoptic segmentation, unifies both semantic and instance-level labeling, offering a comprehensive view of scene composition. As summarized by Minaee et al. (2020), semantic segmentation has been performed using methods such as thresholding (Otsu, 1979), histogram-based bundling, region-growing (Dhanachandra et al., 2015), k-means clustering (Nock and Nielsen, 2004), watershedding (Najman and Schmitt, 1994), to more advanced algorithms such as active contours (Kass et al., 2004), graph cuts (Boykov et al., 2001), conditional and Markov random fields (Plath et al., 2009), and sparsity-based methods (Starck et al., 2005; Minaee and Wang, 2017). While the literature on segmentation is vast and rapidly evolving, a selection of backbone architectures has been particularly influential in shaping current probabilistic segmentation models. The focus here is not on exhaustiveness, but on those models most relevant to the development of uncertainty-aware approaches.
In particular, following the successful application of CNNs (LeCun et al., 1998), image segmentation experienced rapid progress driven by increasingly powerful and specialized deep architectures. Notably, the Fully Convolutional Network (FCN) (Shelhamer et al., 2014) adapted the AlexNet (Krizhevsky et al., 2012), VGG16 (Simonyan and Zisserman, 2014) and GoogLeNet (Szegedy et al., 2014) architectures to enable end-to-end semantic segmentation. Furthermore, other CNN architectures such as DeepLabv3 (Chen et al., 2017), and the MobileNetv3 (Howard et al., 2019) have also been commonly used. As the research progressed, increasing success has been observed with encoder-decoder models (Noh et al., 2015; Badrinarayanan et al., 2015; Yuan et al., 2019; Ronneberger et al., 2015). Initially developed for the medical applications, Ronneberger et al. (2015) introduced the U-Net, which successfully relies on residual connections between the encoding-decoding path, to preserve high-frequency details in the encoded feature maps. To this day, the U-Net is still often utilized as the default backbone model for many semantic segmentation and even general image generation architectures (Ho et al., 2020; Song et al., 2020), particularly in the medical domain. In fact, reports of recent research indicate that the relatively simple U-Net (Isensee et al., 2020) still outperform more contemporary and complex models (Eisenmann et al., 2023; Isensee et al., 2024).
@@ 3 Probabilistic Image Segmentation
Assuming random-variable pairs $(\mathbf{Y},\mathbf{X})\sim P_{\mathbf{Y},\mathbf{X}}$ that take values in $\mathcal{Y}\in\mathbb{Z}^{K\times H\times W}$ and $\mathcal{X}\in\mathbb{R}^{C\times H\times W}$, respectively, then instance $\mathbf{y}$ can be considered as the ground-truth of a $K$-class segmentation task and instance $\mathbf{x}$ as the query image. The variables $H$, $W$ and $C$ correspond to the image height, width and channel depth, respectively. Conforming to the principle of maximum entropy, the optimal parameters given the data (i.e. posterior) subject to the chosen intermediate distributions can be inferred through Bayes Theorem as
@@
assuming independence of $\mathbf{x}$ from $\bm{\theta}$, and where $p(\bm{\theta})$ represents the prior belief on the parameter distribution and $p(\mathbf{y}|\mathbf{x})$ the conditional data likelihood (also commonly referred to as the evidence). After obtaining a posterior with dataset $\mathcal{D}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ containing $N$ images, the predictive distribution for a new datapoint $\mathbf{x}^{*}$ can be denoted as
@@
As evident, both the variability in the empirical data and the inferred parameters of the model influence the predictive distribution. A straightforward approach to quantify the total uncertainty is achieved by obtaining the predictive entropy, $H[\,\mathbf{Y}|\mathbf{x}^{*},\mathcal{D}\,]$, i.e. the entropy of the predictive distribution $p(\mathbf{Y}|\mathbf{x}^{*},\mathcal{D})$, as
@@
where $I$ represents the mutual information. The first term represents the expected information gain about the model parameters $\bm{\theta}$ given the true label $\mathbf{Y}$. A high value indicates the model is very uncertain about its prediction, meaning observing the true label would provide a large information gain. The second term represents the inherent randomness in the data, found by averaging the predictive uncertainty across the posterior of plausible models from Equation (1). This reflects the noise that cannot be reduced, regardless of how much more training data is observed. This distinction allows us to determine if a model’s uncertainty stems from its own ignorance or from ambiguity inherent in the data. Therefore, uncertainty is often split into two types: epistemic uncertainty, which reflects the model’s limited knowledge of its parameters, and aleatoric uncertainty, which captures the irreducible statistical randomness in the data itself. Nonetheless, determining the nature of uncertainty is rarely straightforward. For example, Hüllermeier and Waegeman (2021) stated that “by allowing the learner to change the setting, the distinction between these two types of uncertainty will be somewhat blurred". This sentiment is also shared by Der Kiureghian and Ditlevsen (2009), noting that “in one model an addressed uncertainty may be aleatory, in another model it may be epistemic". Sharing similar views, we highlight the necessity of careful analyses and possible subjective interpretation regarding the topic as we treat the realm of quantifying spatially correlated uncertainty. Especially with increasingly complex methodologies, treating the uncertainties as separate concepts is mostly theoretical (often even more philosophical) and highly non-trivial in practice.
Figure 2: Uncertainty modeling by treating pixel-level confidence values as parameters of a Probability Mass Function.
@@ 3.1 Conventional segmentation
Regardless of the elegantly formulated Bayesian posterior, most practical approaches make use of so-called “deterministic” segmentation networks. Such models are trained by Maximum Likelihood Estimation (MLE), specified as
@@
which simplifies the training procedure by taking a point estimate of the posterior. This approximation improves as the training data increases and the model parameter variances approach zero. As such, MLE does not include any prior knowledge on the structure of the parameter distribution. This, however, can be achieved through Maximum A Posteriori (MAP) estimation with
@@
For example, assuming Gaussian or Laplacian priors leads to regularizing the $L_{2}$ norm (also known as ridge regression or weight decay) or $L_{1}$ norm of $\bm{\theta}$, respectively (Figueiredo, 2001; Kaban, 2007). To model $p(\mathbf{y}|\mathbf{x},\bm{\theta})$, we make use of a function $f_{\bm{\theta}}:\mathbb{R}^{C\times D}\rightarrow\mathbb{R}^{K\times D}$ that infers the parameters of a Probability Mass Function (PMF). For instance, consider spatial image dimensions $D$ and a CNN with $\mathbf{a}=f_{\theta}(\mathbf{x})$. Then, we can write
@@
with channel-wise indexing over the denominator, which is commonly known as the SoftMax activation (Figure 2). Hence, this approach is probabilistic modeling in the technical sense, although it is not referred to as such in common nomenclature. As such, the approximated distribution can represent and localize uncertain regions. However, the implicit pixel-independence assumption
@@
omits information on structural variation in the segmentation masks. In conventional classification, factorizing the categorical distribution is typically regarded as a logical simplification. In probabilistic segmentation however, this assumption has caused the emergence of a distinct research direction (See Figure 3).
Key takeaways: Conventional models yield independent uncertainty estimates for each pixel, which omits spatial coherence. This failure to model spatial correlation precludes realistic inference and inhibits discerning common underlying causes for uncertainty in neighboring regions.
@@
Figure 3: Illustration of the likelihood function in segmentation models. Color intensities reflect the normalized confidence values that can be interpreted as probabilities. (a) The continuous output results in a horizontal gradient. (b) Maximum likelihood thresholding can be applied. (c) However, the coherence of the segmentation suffers when sampling because information on neighboring pixels is disregarded.
3.2 Review organization and structure
The body of this review is structured around three interconnected concepts that serve as the fundamental organizing framework for the field of uncertainty in image segmentation. This structure is intended to facilitate cross-domain synthesis and enable researchers from diverse backgrounds to navigate the literature effectively. Consistent with the decomposition in Equation 2, we structure the methods based on feature-level (Section 4) and parameter-level (Section 5) uncertainty modeling. We then detail the utility of these uncertainties for four key downstream tasks in Section 6, linking them to methods via Tables 14, where the connection to practical application domains is summarized in Table 5. Following this structured overview, we provide a detailed discussion (Section 7) to synthesize key insights, culminating in practical recommendations for the field (Section 8).
Core concepts: Methods: How uncertainty is mathematically modeled and quantified. Tasks: What specific goals are accomplished using these uncertainty estimates. Applications: Where these methods and tasks are applied in real-world domains.